OpenStudy (anonymous):

2x+8y=2 -2x-4y=6 solve by using elimination by addition or subtraction

OpenStudy (hunus):

\[2x+8y=2\] \[-2x-4y=6\] adding the second equation to the first yields \[2x-2x+8y-4y=2+6\] \[4y=8\] \[y=2\] \[2x+8(2)=2\] \[2x=-14\] \[x=-7\] Thus \[y=2\] \[x=-7\]

OpenStudy (anonymous):

where did you get 14?

OpenStudy (hunus):

\[2x+8(2)=2\] \[2x+16=2\] \[2x+16 -16=2 -16\] \[2x+16=2\] \[2x=-14\] \[x=-7\]

OpenStudy (anonymous):

oh i see i didnt take that extra step